OpenStudy (anonymous):

17b + 4 = -4 1/2. Please solve for b.

OpenStudy (anonymous):

b=-1/2 welcome :-]

OpenStudy (anonymous):

17b +4 =-4 1/2 17b = -4 1/2 -4= -8 1/2 b=- 8 1/2/17 b=-1/2

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Can you help again? @HunterAshmender

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Thanks. -2p - 4 =9

OpenStudy (anonymous):

I got 2.1. Is that right?

OpenStudy (anonymous):

p=-13/2

OpenStudy (anonymous):

so thats -6.1

OpenStudy (anonymous):

reduced its 6.5

OpenStudy (anonymous):

Ok thank you again. 6=2x + 3

OpenStudy (anonymous):

x=3/2

OpenStudy (anonymous):

is that one of your options?